TAROT

Télescopes à Action Rapide pour les Objets Transitoires

Status of TAROT observatories

1. Historical dates for cameras and softwares

1998-01-22 : First light of TAROT Calern (Messier 13) with Apogee AM13 camera (Kaf1300)
2001-06-13 : Apogee AM13 camera replaced by Cemes camera (TH7899)
2001-11-16 : Cemes camera replaced by Audine camera (KAF1600)
2002-01 : Audine camera replaced by the Cemes camera (TH7899)
2004-03-08 : Cemes camera replaced by Andor camera (Marconi 42-40)
2005-03-08 : A new version of Majordome is on
2005-09-12 : PC-Cam cable of Andor was replaced & Tarot6 computer replaced
2006-09-15 : First light of TAROT La silla + VPN ESO.
2007-01-30 : new computer tarot10 arrived at Calern
2007-01-30 : TAROT La Silla network is now VPN oamp.fr
2009-01-10 : TAROT La Silla computers 1,2,3,5,6,8 are replaced with Linux + ROS
2009-09-14 : First robotic night for Zadko.
2009-11-17 : TAROT Calern computers 1,5,6 are replaced with ROS
2010-04 : Beginning of robotic observations with Zadko Australia.
2011-11-25 : TAROT La Silla computers 2,4,5,8,9 are replaced by DELL Linux
2012-12-21 : Zadko Australia roof replaced and PLC Burgess installed
2013-11-30 : Zadko Australia Strehl computer replaced and Andor camera replaced by USB Andor.

2. Satellites devoted to GRB prompt emission in high energy

2000-08 : BATSE end mission
2000-10-10 : HETE launched
2004-nov : SWIFT notices
2009-jan : Antares notices
2009-fev : Fermi notices
2009-oct : Loocup notices

3. Observatory reliabilities

3.1. TAROT Calern

257.5 days of technical problems from 2004-03-16 to 2013-01-06 (3395) => 92 % of reliability

2002-08-10 to 2002-12 : a strong storm destroyed many things in the observatory (4 months)
2003-12-31 to 2004-01-03 : bug in the new driving software (3 days)
2004-02-21 to 2004-03-16 : Cemes camera failed and was replaced by the Andor (1 month)
2004-05-07 to 2004-05-10 : Link between two computers failed (3 days).
2004-06-22 to 2004-06-23 : Link between meteo-station and Tarot was broken (1 day).
2004-06-28 to 2004-07-11 : Link between meteo-station and Tarot was broken (14 days).
2004-10-01 to 2004-10-05 : Filter wheel blocked (5 days).
2004-10-23 to 2004-10-24 : A bug crashed the Majordome soft (1 day)
2004-12-28 to 2005-01-03 : The roof mechanics failed (7 days)
2005-03-16 to 2005-03-16 : Hour axis blocked (1 day)
2005-04-08 to 2005-04-10 : Hour axis blocked (2 days)
2005-05-06 to 2005-05-10 : Network failed (OCA) and hour axis blocked (3 days)
2005-05-14 to 2005-05-16 : Tarot6 computer crashed (due to a storm ?) (3 days)
2005-05-28 to 2005-05-30 : Tarot6 computer crashed (3 days)
2005-06-12 to 2005-06-12 : Tarot6 computer crashed (1 day)
2005-06-29 to 2005-06-30 : Tarot6 computer crashed (0.5 day)
2005-07-03 to 2005-07-04 : Tarot6 computer crashed (1 day)
2005-08-13 to 2005-08-17 : Tarot6 computer crashed (3 day) + Tarot2 Majordome software crashed.
2005-08-24 to 2005-08-24 : Tarot6 computer crashed (1 day)
2005-08-28 to 2005-08-28 : Tarot6 computer crashed (1 day)
2005-09-12 to 2005-09-16 : Tarot6 computer replaced, optic cleaned and aligned (5 days)
2005-10-28 to 2005-11-02 : Socreq software crashed during the hollidays ! (5 days)
2006-01-29 to 2006-02-10 : Mecanical link of the delta axis was broken (13 days)
2006-03-19 to 2006-03-23 : Tarot6 computer memory failed but no incidence on observations (0 days) => T6 replaced by the "mulet"
2006-04-21 to 2006-04-21 : Tarot6 computer disk full (1 day)
2006-05-04 to 2006-05-04 : Bad init on the delta axis (0.5 day)
2006-05-11 to 2006-05-11 : A new set of filters have been installed (0.5 day)
2006-05-21 to 2006-05-22 : Hacker intrusion in Tarot6 computer (1 day)
2006-05-21 to 2006-05-22 : Hacker intrusion in Tarot6 computer (1 day)
2006-05-29 to 2006-05-29 : Tarot6 "mulet" computer is converted into Tarot9 and the new Tarot6 is on (0 day)
2006-05-31 to 2006-06-02 : A bad filter focus configuration leads to unfocussed images (1.5 day)
2006-08-01 to 2006-08-02 : The red button was not reactived after an intervention (1 day)
2006-09-16 to 2006-09-18 : Electrical problem due to a storm (3 days)
2006-10-10 to 2006-10-13 : Bad reinit of the Ra-Dec motors (3 days)
2006-11-14 to 2006-11-15 : Bad reinit of the Ra-Dec motors (0.5 days)
2006-11-17 to 2006-11-20 : Cleaning of cog wheels. (0 days)
2006-12-05 to 2006-12-08 : Aluminium coating at Nice observatory + focus setting (3 days)
2007-01-03 to 2007-01-03 : Bad reinit of the Ra-Dec motors (0.25 days)
2007-01-03 to 2007-01-04 : Hour angle Switch blocked (0.70 days)
2007-02-01 to 2007-02-01 : Bad reinit of the Ra-Dec motors (0.5 days)
2007-02-12 to 2007-02-12 : GPS time installed (0 day)
2007-03-07 to 2007-03-08 : Bad status returned by switch stops of focus (2 days)
2007-06-12 to 2007-06-14 : Camera cable unpluged (2 days)
2007-05-15 to 2007-07-10 : GPS time was out on 15th May and then replaced 10th July (0 days)
2007-07-03 to 2007-07-10 : Hour angle screw not enough strong, trailed images (0 days)
2007-07-22 to 2007-07-24 : UPS anomaly ? (2 days)
2007-08-23 to 2007-08-24 : Communication problem between Majordome2 and Socreq (1 day)
2007-09-02 to 2007-09-04 : Storm protector default (3 days)
2007-11-05 to 2007-11-06 : Unidentified electrical problem (1 day)
2007-11-26 to 2007-11-26 : Image disk full (0 day)
2007-11-07 to 2007-11-06 : Electrical problem due to wind and EDF (3 days)
2007-12-22 to 2008-01-05 : Power supply of tarot1 failed (14 days)
2008-01-15 to 2008-01-16 : Roof mechanical problem (2 days)
2008-02-07 to 2008-02-07 : Internet connexion problem outside Calern (0 day)
2008-06-30 to 2008-07-07 : Power supply failure for telescope controlers (6 days)
2008-09-16 to 2008-09-19 : Telescope stopped on the western sensor (3 days)
2008-11-01 to 2008-11-14 : Camera.exe failure (14 days)
2009-02-12 to 2009-02-12 : Tarot6 seems to be frozen. Rebooted (0 day)
2009-09-16 to 2009-11-20 : Tarot1 disk is dead. Great maintenance (65 days) [FV, S, ML,AK]
2010-01-25 to 2010-01-28 : Maintenance, UPS (0 day) [AK, ML]
2010-02-11 to 2010-02-11 : Too much snow for the roof -> Manual (0 day) [SR]
2010-02-27 to 2010-03-04 : Mecanical link of the alpha axis was broken (5 days)
2010-05-15 to 2010-05-20 : Bug of GOTO telescope + init (5 days) [ML, AK, SR]
2010-06-14 to 2010-06-18 : Power supply 24V of motors is dead. Replaced. (5 days) [SR]
2010-09-10 to 2010-09-10 : Maintenance axis delta (0 day) [SR]
2010-09-26 to 2010-09-26 : Defocused images (0.5 day) [SR]
2010-10-01 to 2010-10-01 : Telescope.exe crashed (1 day) [SR]
2010-10-21 to 2010-10-23 : Maintenance, mechanics, electric (2 days) [OHPTec, YR, XR, RS, AK, SR]
2010-11-01 to 2010-11-02 : Socreq no response to Majordome (2 days) [AK]
2011-01-13 to 2011-01-13 : Telescope out of stops (1 days) [SR]
2011-01-24 to 2011-01-24 : Flat lamp replaced (0 day) [SR]
2011-02-26 to 2011-03-04 : Telescope out of stops due to software update (6 days) [AK]
2011-03-26 to 2011-04-04 : Declination motor dead. Replaced. (6 days) [SR]
2011-11-17 to 2011-11-17 : Focusing model measurements. (0,5 day) [AK]
2011-11-17 to 2011-11-18 : Meteo cloud sensor installation. (0 day) [SR+John Moore]
2011-12-07 to 2011-12-07 : Tests of hardness for meteo sensor connection. (0,5 day) [AK]
2011-12-14 to 2011-12-14 : Roof remains opened in the morning. (0 day) [SR]
2012-01-17 to 2012-01-18 : Database maintenance (1 days) [AK, FV]
2012-02-04 to 2012-02-04 : Slow motion problems in hour angle (0.5 day) [AK]
2012-02-08 to 2012-02-08 : Bugs in the new version of the make_scenes in Majordome.exe (0.5 day) [AK]
2012-02-11 to 2012-02-11 : Motion problems in hour angle. Cable connectors repluged  (1 day) [SR]
2012-03-04 to 2012-03-04 : Software updates and tests in the beginning of the night (0.5 day) [AK]
2012-03-08 to 2012-03-08 : Software updates and debug in the beginning of the night (0.5 day) [AK]
2012-03-29 to 2012-03-29 : Roof stop sensor blocked (0.5 day) [SR]
2012-05-11 to 2012-05-12 : Power supply off because a strong storm. Tarot1 switched on manually (0.5 day) [SR]
2012-08-13 to 2012-08-13 : Electricity shut down of CADOR. Only local planning (1 day) [Julien Lecubin]
2012-08-26 to 2012-08-26 : Electricity shut down of CADOR due a storm (1 day) [itself]
2012-08-30 to 2012-09-16 : Flat field lamp is broken. Use of older flats (0 day) [SR]
2012-09-26 to 2012-09-26 : Flat field lamp is replaced (0 day) [SR]
2012-09-30 to 2012-10-02 : CADOR link broken due to the file server Aldebaran at OHP (3 day) [JL]
2012-10-16 to 2012-10-18 : CADOR link broken due to the file server Aldebaran at OHP (2 day) [JL]
2012-10-19 to 2012-10-22 : Virus detection by OCA that broke the inthernet link (3 day) [Christophe Limonta]
2012-11-05 to 2012-11-07 : Electricity + UPS problem + pb switch of Tarot1  (1.5 day) [SR]
2012-11-29 to 2012-11-30 : Electricity + UPS problem + pb switch of Tarot1  (1 day) [SR+ workshop]
2012-12-04 to 2012-12-11 : Coating of the M1 and M2 mirrors  (7 days) [SR]
2013-02-09 to 2013-02-11 : CADOR link broken due to the file server Aldebaran at OHP (3 day) [JL,GC]
2013-02-11 to 2013-02-13 : CADOR down at OHP (2 day) [JL]
2013-06-08 to 2013-06-09 : CADOR link broken due to the file server Aldebaran at OHP (2 day) [JL,AK]
2013-08-22 to 2013-09-02 : CADOR svn checkout failed config.ini files (12.5 day) [SR, AK]
2013-12-18 to 2013-12-27 : Tarot9 MySQL serveur disk full cause to the daily logs (9 days) [SR, AK]
2013-12-22 to 2013-12-22 : GPS date symmetricom is out of date but corrected by the path (0 days) [SR, AK]
2013-12-25 to 2014-01-08 : CADOR Sargas switched out (0 days) [Julien charpin]
2014-01-17 to 2014-01-17 : Maintenance web site problem solved with a big SVN update (0 days) [SR, AK]
2014-02-28 to 2014-02-28 : Problem with a cable of the roof due to 10cm snow (1 day) [SR]
2014-04-04 to 2014-04-04 : Ethernet switch blocked only for tarot1 computer (0.5 day) [SR]
2014-06-04 to 2014-06-06 : No internet access (0 day) [PR, AK]
2014-08-03 to 2014-10-08 : Reboot of tarot6 and tarot10 due to pb of shared disks (0 days) [AK, PR]
2015-01-23 to 2015-01-24 : Majordome crash (Tcl 8.6) (1 day) [AK, SR]
2015-01-29 to 2015-01-29 : Gardien crash (Tcl 8.6) (1 day) [AK, SR]
2015-02-04 to 2015-03-15 : Tarot10 HDD crash (0 day) [SR, AK,]
2015-02-14 to 2015-03-29 : Roof steel cable broken + TAROT10->TAROT6 [SR, atelier Calern]
2015-06-25 to 2015-06-25 : New PC for CADOR [Julien Charpin]

3.2 TAROT LaSilla

686.5 days of technical problems from 2006-09-21 to 2013-07-02 (2476) => 72 % of reliability

2006-10-10 to 2006-10-12 : Switch problem in the northern roof (2.5 days)
2006-11-16 to 2006-11-22 : VPN break down + socket problem between Majordome2 and socreq (6 days)
2006-11-28 to 2006-12-02 : Socreq crash (3 days)
2007-01-03 to 2007-01-06 : Camera.exe crash (4 days)
2007-01-30 to 2007-01-30 : Rain detector failed (0.5 days)
2007-02-02 to 2007-02-02 : Socreq crach (1 day)
2007-03-10 to 2007-03-12 : Socreq crach (1.5 day) 1 alert lost
2007-03-12 to 2007-03-13 : Socreq crach (0.5 day)
2007-04-10 to 2007-04-13 : Camera.exe crach (3 days)
2007-05-09 to 2007-05-09 : Socreq.exe crach (1 day) 1 alert lost
2007-05-19 to 2007-05-20 : Socreq.exe crach (1.5 day)
2007-05-21 to 2007-05-21 : Socreq.exe crach (0.5 day) + error supervision (0.5 day)
2007-05-22 to 2007-06-10 : Motor power supply crach + maintenance + GPS time installed (19 day)
2007-06-11 to 2007-06-15 : Maintenance (4 day)
2007-06-22 to 2007-06-23 : Motor focus switch stop (1.5 day)
2007-07-09 to 2007-07-10 : Socreq.exe crach (1.5 day)
2007-07-20 to 2007-07-20 : Patch in trireq to prevent Socreq.exe crach (0 day)
2007-07-20 to 2007-08-05 : Motor power supply crach (16 days)
2007-08-05 to 2007-08-15 : Maintenance (0 day)
2007-08-20 to 2007-08-21 : Motor power supply crach (2 days)
2007-08-25 to 2007-08-26 : Bad manual launch of Trireq.exe (1 day)
2007-10-12 to 2007-10-16 : Camera.exe crach due to socket connection (4 days)
2007-10-16 to 2007-10-27 : Cable of z+ sensor unpluged to the declination axis socket (11 days)
2007-10-27 to 2007-10-28 : Camera.exe crach due to MySQL socket connection (2 days)
2007-10-28 to 2007-10-31 : Computer tarotchili8 switched off (4 days)
2007-11-16 to 2007-11-17 : Air cooling set 1 in the shelter cannot start (1 day)
2007-11-29 to 2007-11-29 : Filter motor cannot move. Telescope and axes relays off/on to reboot (1 day)
2007-12-12 to 2007-12-12 : Computers tarotchili1,3,5,8 switched off (1 day)
2008-01-01 to 2008-01-03 : Bug in requeteur.exe (0 day)
2008-01-07 to 2008-01-09 : Computer tarotchili8 switched off (1 days)
2008-01-09 to 2008-01-16 : VPN Router crashed (7 days)
2008-02-21 to 2008-04-13 : VPN Router crashed (51 days)
2008-03-03 to 2008-04-13 : Switch of the motor power supply crashed
2008-04-13 to 2008-04-29 : Switch of the motor power supply crashed (16 days)
2008-06-19 to 2008-06-21 : Maintenance during bad weather (3 days)
2008-06-28 to 2008-07-02 : Power supply general of la Silla failure (4 days)
2008-07-05 to 2008-07-06 : Tarotchili8 failure. Socreq failure (0.5 day)
2008-07-08 to 2008-08-04 : Wifi deconnection + Power supply general of la Silla failure (24 days)
2008-08-16 to 2008-08-29 : Power supply general of la Silla failure (14 days)
2008-08-29 to 2008-11-03 : Router VPN config problem + tarotchili4 power supply failure (67 days)
2008-11-24 to 2008-11-26 : La Silla power supply failure (2 days)
2008-11-27 to 2008-11-30 : Camera.exe crashed because Tarotchili5 remains off (3 days)
2008-12-09 to 2008-12-23 : Wifi deconnection (14 days)
2009-01-02 to 2009-01-16 : Maintenance. Change of Dec. motor + change computers 1,2,3,5,6,8 (14 days)
2009-01-17 to 2009-01-18 : Tarotchili-4 frozen (2 days) [VM=repaired by Vincent Megevan]
2009-01-31 to 2009-02-01 : Tarotchili-1 frozen. Rebooted by blackbox (1 day)
2009-02-06 to 2009-02-11 : Tarotchili-4 frozen (5 days) [VM]
2009-02-11 to 2009-02-16 : Tarotchili-4 frozen + global power supply down (5 days) [VM]
2009-02-16 to 2009-02-16 : Tarotchili-5 frozen. Rebooted by blackbox (0.5 day)
2009-02-24 to 2009-03-01 : Tarotchili-4 shut down. (5 days) [ChTec=repaired by chilean technicians]
2009-03-01 to 2009-03-05 : Roof on manual mode. (4 days) [ChTec=repaired by chilean technicians]
2009-04-04 to 2009-04-07 : Roof on manual mode. (3 days) [ChTec].
2009-04-09 to 2009-04-11 : Bug in Majordome due to the fix of the roof openning at noon. (2 days) 
2009-04-29 to 2009-05-16 : Electric shutdown at 13h28 TU (~13h00 for camera) (17 days)  [VM]
2009-05-25 to 2009-05-26 : Roof problem during the night. (1 day) [ChTec].
2009-06-01 to 2009-06-16 : Electric shutdown at 13h50 TU (~13h27 for camera) (17 days)
2009-07-01 to 2009-07-10 : Electric shutdown for ESO maintenance (10 days)  [ChTec]
2009-07-11 to 2009-07-11 : Electricity switched on [ChTec]
2009-07-11 to 2009-08-15 : Tarotchili-4 not switched on and everybody on hollidays (35 days)
2009-08-15 to 2009-08-15 : Tarotchili-4 switched on [VM]
2009-08-15 to 2009-09-15 : Roof electric bloc jamed (30 days) [ChTec].
2009-10-08 to 2009-10-11 : Roof electric bloc jamed at 12h01 UT (4 days) [ChTec].
2009-10-11 to 2009-10-13 : CADOR BDD connexion problem  (2 days).
2009-10-19 to 2009-10-21 : Tarotchili5 and tarotchili8 switched off (3 days) [VM].
2009-10-29 to 2009-11-02 : Tarotchili5 switched off (3 days) [ChTec].
2009-11-06 to 2009-11-12 : Electric shutdown from ESO (6 days) [VM].
2009-11-17 to 2009-11-28 : Electric shutdown from ESO (1 day) [Alexandre Santerne].
2010-01-03 to 2010-01-03 : Manual switch off during day remote maintenance (1 day) [ChTec].
2010-01-07 to 2010-01-13 : Electric shutdown from ESO (5 days) [ChTec].
2010-02-23 to 2010-02-24 : Electric shutdown from ESO (2 days) [ChTec].
2010-02-27 to 2010-03-04 : Electric shutdown due to the great earthcake (6 days) [ChTec].
2010-03-04 to 2010-03-16 : Breakers problems for AE02 electrical power (12 days) [OHPTec, YR, ALVS, XR].
2010-03-16 to 2010-03-28 : Maintenance visit, mechanics, electricity (12 days) [OHPTec, YR, ALVS, XR].
2010-04-16 to 2010-04-23 : Tarotchili5 switched off (7 days) [ChTec].
2010-05-01 to 2010-05-11 : Tarotchili5 switched off (11 days) [IMCCETec, FV, JB].
2010-05-11 to 2010-05-11 : Tarotchili1 takes relay of Tarotchili5 (0 days) [AK].
2010-05-12 to 2010-05-21 : Maintenance visit, computer science (10 days) [IMCCETec, FV, JB].
2010-06-23 to 2010-06-28 : Tarotchili4 camera.exe, VNC, cgi-apache crashed (5 days) [VM]
2010-07-27 to 2010-08-05 : Tarotchili4 pb camera non coolled (8 days) [AK]
2010-08-02 to 2010-08-04 : Tarotchili1 pb compils after cvs update (2 days) [AK,ML]
2010-09-27 to 2010-10-08 : Tarotchili4 power supply breakdown. Replaced (12 days) [ChTec]
2010-10-08 to 2010-10-10 : Bug in Grenouille produces a delay for image processing (0 days) [AK]
2011-01-03 to 2011-01-03 : Pulse problem of the rain detector seen vizualy (0 days) [AK]
2011-02-27 to 2011-03-03 : Emergency button pushed due to the pulse problem of the rain detector (5 days) [Gregory Lambert]
2011-03-06 to 2011-03-14 : Emergency button pushed due to the pulse problem of the rain detector seen vizualy (8 days) [GL]
2011-03-16 to 2011-04-01 : Emergency button pushed due to the pulse problem of the rain detector + solved (16 days) [ChTec+XR]
2011-06-06 to 2011-06-09 : Emergency button pushed to prevent snowing weather without rain detector (3 days) [ChTec+GL]
2011-06-19 to 2011-06-19 : Rain announced without rain detector. Manual mode actived during the alert (0 days) [ALVS]
2011-07-08 to 2011-07-22 : Rain in the telescope without rain detector. Manual breakers switched (14 days) [ChTec+XR+ALVS+AK]
2011-07-29 to 2011-08-01 : Snow announced without rain detector. Manual mode actived during the alert (0 days) [AS+AK]
2011-10-17 to 2011-10-24 : Telescope crash (EAlimAxes=0) (8 days) [AK]
2011-10-25 to 2011-10-25 : VPN IP change (0 day) [ChTec + ALVS + oamp.frTechs ]
2011-10-30 to 2011-11-02 : Tarotchili4 power supply breakdown. Replaced (4 days) [GL]
2011-11-03 to 2011-11-08 : Tarotchili4 power supply breakdown. Replaced (5 days) [GL]
2011-11-22 to 2011-12-05 : Maintenance. Computer Tarotchili4 replaced, new ones installed. Optics cleaned (14 days) [YR+XR+FD]
2011-12-21 to 2011-12-21 : Telescope hardware crash (sensor stop reached at the beginning of the night) (0.5 days) [AK]
2012-01-13 to 2012-01-22 : Telescope maintenance, computer changes (multiple stops) (9 days) [AK, FV, JB, PT]
2012-01-29 to 2012-01-30 : Telescope coordinate system has changed (1 day) [AK]
2012-03-09 to 2012-03-09 : Majordome software bug (1 day) [AK]
2012-04-16 to 2012-04-20 : Cable of z+ sensor unpluged to the declination axis socket (4 day) [XR, GL]
2012-08-13 to 2012-08-13 : Electricity shut down of CADOR. Only local planning (1 day) [Julien Lecubin]
2012-08-14 to 2012-08-17 : Problem with MySQL connection from CADOR. Only local planning (3 day) [Julien Lecubin]
2012-09-30 to 2012-10-02 : CADOR link broken due to the file server Aldebaran at OHP (3 day) [JL]
2012-10-16 to 2012-10-18 : CADOR link broken due to the file server Aldebaran at OHP (2 day) [JL]
2013-02-08 to 2013-02-11 : CADOR ankaa computer problem and no replica for planning (3 day) []
2013-02-09 to 2013-02-11 : CADOR link broken due to the file server Aldebaran at OHP (3 day) [JL,GC]
2013-02-11 to 2013-02-11 : Telescope blocked due to electrical switch AXES=0 (0.5 day) [AK]
2013-02-11 to 2013-02-13 : CADOR down at OHP (1.5 day) [JL]
2013-04-13 to 2013-04-15 : Ethernet switch of the AE03 is dead and replaced (1.5 day) [GL, AK, XR]
2013-04-19 to 2013-04-22 : The replaced ethernet switch of the AE03 is blocked. Reboot (4 days) [GL, AK]
2013-05-10 to 2013-05-13 : Telescope on western stop sensor + EAlimAxes=0 (3 days) [AK]
2013-05-28 to 2013-05-28 : Telescope on western stop sensor (1 day) [AK]
2013-06-08 to 2013-06-09 : CADOR link broken due to the file server Aldebaran at OHP (2 day) [JL,AK]
2013-06-30 to 2013-07-01 : The replaced ethernet switch of the AE03 is blocked. Reboot 30T11:39 -> 01T19:30 (1.5 days) [GL, AK, F. Bouchy, Isabelle Boisse, XR]
2013-07-01 to 2013-07-03 : Telescope lost connection with DeltaTau (2 days) [AK]
2013-07-18 to 2013-07-18 : The replaced ethernet switch of the AE03 is blocked. Reboot (1 days) [GL, AK]
2013-07-21 to 2013-07-31 : VPN box of the shelter is dead. Roof opened. Replaced and switch reboot (10 days) [GL, ALVS, ChTec]
2013-08-02 to 2013-08-02 : Ethernet switch of the AE03 is blocked. Reboot (1 day) [GL, AK]
2013-08-06 to 2013-09-18 : Ethernet switch of the AE03 is blocked. System stopped (43 day) [XR, AK]
2013-09-17 to 2013-10-03 : Maintenance (0 day) [XR]
2013-10-09 to 2013-10-11 : Telescope software connection pb with Umac (2 day) [AK]
2013-12-25 to 2014-01-08 : CADOR Sargas switched out (0 days) [Julien charpin]
2014-02-14 to 2014-03-01 : Network change at Marseille, TAROT router not well configured (15 days) [ALVS, ChTec]
2014-06-03 to 2014-06-17 : Switch roof north problem + Mysql saturation of .bin archives (14 days) [AK]
2014-06-19 to 2014-06-29 : Perturbations caused by full disks on tarotchili5 and tarotchili8 (3 days) [AK]
2014-06-27 to 2014-06-29 : camera.exe pb of allocation memory on tarotchili4 (2 days) [AK]
2014-08-02 to 2014-10-12 : DNS not accessible causes bad update of NTP time (0 day) [JC, AK]
2015-02-05 to 2015-03-11 : Crash HDD tarotchili4 camera (34 days) [AK, XR, VM]
2015-02-11 to 2015-03-22 : Maintenance + tarotchili4 + change the focuser (11 days) [XR, YR, RS, AK]
2015-03-23 to 2015-03-30 : Moxa 192.168.10.58 bad init (8 days) [XR, AK]
2015-03-30 to 2015-05-08 : Multiple problems NTP, Camera crashes, CADOR (87 days) [AK, FV]
2015-05-08 to 2015-06-25 : Crash for CADOR [Julien Charpin]
2015-06-25 to 2015-08-16 : New PC for CADOR but configuration problems [Julien Charpin]
2015-08-08 to 2015-08-19 : Storm and roof stayed opened. AE02 in MAN mode [Tech Ch, VM]

3.3. Zadko Australia

176 days of technical problems from 2012-12-21 to 2013-07-02 (345) => 51 % of reliability

2012-05-30 to 2012-12-21 : Dome replacement by a roof (6.5 months) [JM+XR+FD+MB+PT]
2013-01-01 to 2013-01-04 : Flush hosts problem in the Strehl MySQL serveur (4 days) [MB, AK]
2013-01-08 to 2013-01-09 : Flush hosts problem in the Strehl MySQL serveur (1.5 days) [AK]
2013-02-07 to 2013-02-11 : TCS clock advanced by 4h30 / PC UTC (3 days) [AK]
2013-02-07 to 2013-02-07 : M1& M2 cleaned + temporary baffle (0 days) [JM]
2013-02-08 to 2013-02-11 : CADOR ankaa computer send bad datadue to Aldebaran file server (3 day) [JL, GC]
2013-02-11 to 2013-02-13 : CADOR down at OHP (2 day) [JL]
2013-02-14 to 2013-02-14 : DFM blocked on an elevation stop ensor for asteroid 2012DA14 (0.5 day) [AK]
2013-02-14 to 2013-02-16 : DFM focusing change from 3153 to 3184 (1.5 day) [AK]
2013-02-22 to 2013-02-26 : Power supply off due to forgot to turn the power back on for the PLC box (4 days) [JM]
2013-02-26 to 2013-02-26 : DFM focusing change from 3184 to 3171 (0.5 day) [AK]
2013-02-27 to 2013-02-28 : Filter wheel blocked (2 day) [AK]
2013-03-01 to 2013-03-01 : Flush hosts problem in the Strehl MySQL serveur (0 day) [AK]
2013-03-06 to 2013-03-06 : Burgess let in MANUAL position (1 day) [XR]
2013-04-19 to 2013-04-24 : Internet connection is off + flush hosts error (5 days) [AK]
2013-04-26 to 2013-04-26 : TCS clock advanced by 22.6h / PC UTC (1 day) [EC, AK]
2013-05-13 to 2013-05-13 : Internet connection is off + flush hosts error (1 day) [EC, AK]
2013-05-17 to 2013-06-05 : Roof wheels crash (19 days) [JM]
2013-06-13 to 2013-06-14 : TCS clock advanced by 5.3h / PC UTC (1 day) [EC, AK]
2013-06-26 to 2013-06-27 : Electric crash + TCS bad automatic reboot (1 day) [EC, AK]
2013-06-28 to 2013-06-30 : DFM focusing change from 3171 to 3149 + closing roof pb (2 days) [EC, AK]
2013-06-30 to 2013-07-02 : Internet connection is off + flush hosts error (3 days) [EC, AK]
2013-08-03 to 2013-12-01 : Strehl computer is dead (120 days) [DC, EC, AK, MB, JM, Julien charpin]
2014-01-06 to 2014-01-10 : Telescope.exe killed probably by a virus (5 days) [Mark Dickinson, JM, MB]
2014-01-14 to 2014-01-23 : Telescope.exe killed probably by a virus (9 days) [Mark Dickinson, JM, XR]
2014-02-11 to 2014-02-28 : Zadko6 causes a shutdown of the internet rate (17 days) [Mark Dickinson, Katrina Chapman OceanBroadBand]
2014-03-25 to 2014-03-25 : Telescope stopped in the western horizon. Solved with the pad of TCS (0 days) [PT, AK]
2014-04-02 to 2014-06-19 : Roof motor + variator problem + internet interruptions (78 days) [MB, MD]
2014-06-19 to 2014-06-28 : Focusing difficulties + internet interruptions (9 days) [AK]
2014-07-19 to 2014-09-19 : Power supply (UPS ?) off for the blue box DFM (61 days) [AK, JM, MD]
2014-10-05 to 2014-10-06 : Power supply (UPS ?) off for the blue box DFM (2 days) [AK, JM, MD]
2014-10-07 to 2014-10-07 :  DNS have been changed without information  [AK, OceanBroadband]
2014-10-07 to 2014-10-07 :  NTP drift corrected after the DNS correction (0 days) [AK]
2014-10-10 to 2014-10-12 :  Burn off of the bush (4 days) [MB information]
2014-10-13 to 2014-10-13 : UPS upgrade (1 day) [JM]
2015-02-09 to 2015-02-20 : Maintenance ( day) [AK, PT, DT, AdK, JM, DC]
2015-02-22 to 2015-02-22 : DFM breaker (1 day) [Mark Dikinson, AK]
2015-03-12 to 2015-03-20 : Internet connection oceanbroadband broken due to a strom (8 days) [Mark Dikinson, OBB, AK]
2015-03-20 to 2015-0?-?? : Camera no response (?? days) [Mark Dikinson, AK]